# P3-LLM: An Integrated NPU-PIM Accelerator for LLM Inference Using Hybrid Numerical Formats

### v1

@@ IV-B Attention-Score Quantization
In order to fully exploit the hardware efficiency of KV-cache quantization, it is essential to quantize additional operands, e.g., attention-scores, within the self-attention module. Although attention-scores have little impact on the overall memory footprint (see Section III-A), they play a critical role in the computation flow during self-attention. Consider a low-precision PCU supporting 8-bit multiplication, if the attention scores remain in FP16, the PCU cannot be fully utilized to accelerate self-attention. Instead, the quantized value cache must be transferred to NPU and rely on the FP16 compute units to perform multiplication with attention-scores, thus diminishing the advantages of high internal PIM bandwidth for LLM decoding. Unfortunately, existing KV-cache quantization approaches, such as QServe [qserve] and Oaken [oaken], typically keep attention-scores in FP16, which hinders the usage of low-precision hardware for multiplication with the value cache. A straightforward solution is to quantize attention-scores with INT8 [sage-attn], but this introduces two drawbacks. First, INT8 attention-scores cause noticeable degradation in perplexity as shown in Table II. Second, as we will demonstrate in Section IV-D, INT8 can lead to significant accuracy loss for activation quantization. Since attention-scores behave the same as activations during the MAC operation on hardware, it is desirable to identify a good numerical format that balances quantization accuracy and hardware complexity.
To address this, we propose an unsigned 8-bit floating-point format, FP8-S0E4M4, that contains a 4-bit mantissa and a 4-bit exponent with an exponent bias of 15-15. There are two insights that drive our design choice. First, because the attention-scores are produced after a softmax function (see Section II-A), its numerical range always lies between 0 and $1$, eliminating the need for a sign bit in its encoding. Second, recall FP16 has a 5-bit exponent with a bias of 15-15, which provides an exponent range of [14,15][-14,15]. However, since attention-scores are always less than 1, there is no need to use positive exponent values, leading to an effective exponent range of [14,1][-14,-1] with 14 distinct values. Thus, a 4-bit exponent is sufficient to represent the attention-score range, which leaves 4 bits to store mantissa with better numerical fidelity. Table II reflects that FP8-S0E4M4 outperforms both INT8 and FP8-E4M3 for attention-score quantization, and achieves near-lossless model performance.
@@ IV-C Weight Quantization
For weight quantization, we build upon an existing numerical format, BitMoD [bitmod], which adaptively remaps the redundant negative zero encoding of FP4 to some pre-defined special values. On top of the basic FP4 quantization values {± $\{\pm\,0,\,\pm\,0.5,\,\pm\,1,\,\pm\,1.5,\,\pm 2,\,\pm 3,\,\pm 4,\,\pm 6\}$, BitMoD introduces four additional special values {± $\{\pm\,5,\,\pm\,8\}$, where one of them can be selected to replace the negative zero. By searching for the optimal special value for each group of weights, it reduces the quantization error compared to asymmetric integer quantization with minimal hardware overhead. Notably, since weights and KV-cache are mapped to the same operand on the MAC hardware, a decoder is required to accommodate the data types of both weights and KV-cache (i.e., BitMoD and INT4-Asym), which we discuss in Section V-C.
@@ V P3-LLM Architecture
This section outlines the architectural details of P3-LLM, which features: (a) an efficient operator mapping that allows most memory-intensive operators to be mapped to PIM during decoding, (b) a quantized matrix multiplication dataflow that applies operator fusion to minimize the runtime overhead of quantization and dequantization, (c) a high-throughput and area-efficient PCU supporting hybrid numerical formats.
V-A PIM Offloading Opportunities with P3-LLM
@@
The proposed W4A8KV4P8 quantization allows memory-bound matrix multiplication to be executed efficiently on low-precision hardware, as illustrated in Fig. 6(a). During the decoding stage, the host NPU performs all element-wise operations such as RoPE and softmax in high precision (FP16 or FP32), while most GEMM/GEMV operators such as linear layers are offloaded to the low-precision PCU. One exception is $Q\cdot K^{\text{T}}$}, whose mapping depends on whether the key cache is quantized pre-RoPE or post-RoPE. In the former case, the quantized key cache lacks positional information during decoding, which can produce incorrect outputs if multiplied directly with the query. Therefore, an online RoPE operation must be applied to the key cache in each decoding iteration. Fortunately, RoPE is essentially an element-wise operation with negligible runtime overhead[flashinfer]. Since RoPE is performed on NPU and its outputs are high-precision, we opt to calculate $Q\cdot K^{\text{T}}$} on NPU if the key cache is quantized before RoPE. In contrast, the post-RoPE-quantized key cache can be directly multiplied with the query, which permits $Q\cdot K^{\text{T}}$} to be computed on the low-precision PCU.
V-B Optimization on Quantized Dataflow
@@
Quantization Granularity. The quantization granularity has a direct impact on the computation flow. For query and KV-cache, we employ per-head quantization, where a group of $D_{h}$ elements are quantized together and $D_{h}$ is the head size. This allows immediate quantization of each newly generated token and eliminates the need to buffer a window of FP16 KV-cache as done in prior works [mant, kivi]. Additionally, weights are quantized per-group with a group size of $128$ following BitMoD [bitmod]. Activations are quantized per-token using FP8-E4M3, as described in Section IV-D. Finally, attention-score quantization does not require scaling since FP8-S0E4M4 already covers the required numerical range. Instead, we directly round the least significant bits of the FP16 attention-scores and keep their highest 4 mantissa bits.
@@
Quantized GEMV Dataflow. As shown in Fig. 6(b), P3-LLM implements operator fusion to minimize the runtime overhead of (de)quantization, which needs to be executed on NPU using high-precision arithmetic. For linear layers, the dequantization scaling is performed after the completion of matrix multiplication. For $Q\cdot K^{\text{T}}$}, since the post-RoPE key cache contains per-channel smoothing factors ($S_{\text{SF}}$), we fuse $S_{\text{SF}}$ into query via element-wise multiplication prior to FP8 quantization. Similarly, for $P\cdot V$}, we fuse the per-value-head scaling factors ($S_{V}$) into attention-scores before FP8 quantization. However, a potential issue of fusing $S_{V}$ is that it may cause the resulting attention-scores’ magnitude to become larger than one, which is outside the range of FP8-S0E4M4. To prevent this, we apply a second-level scaling that divides $S_{V}$ with its maximum, resulting in $S_{V}\in[0,1]$}. Note that the quantization scaling factor is always non-negative by definition in Eq. 1, so fusing $S_{V}$ still produces unsigned attention-scores. This second-level scaling factor is later multiplied back into the result of $P\cdot V$}.
@@
V-C P3-LLM Microarchitecture
To support the proposed W4A8KV4P8 scheme with hybrid numerical formats in PIM, we design an efficient PCU on top of HBM-PIM [samsung-pim]. Fig. 6(c) left shows the P3-LLM architecture containing a host NPU and an HBM-PIM. One PIM channel contains 8 PCUs, each of them is shared by 2 banks to amortize the area overhead. During the decoding stage, the NPU sends activations and attention-scores to the input register of PCU, while weights and KV-cache are read from DRAM banks to perform MAC operations. Given that weights and KV-cache are quantized to 4 bits, we choose to equip each PCU with 64 multipliers to match the 256-bit column access granularity of HBM, effectively quadrupling the compute throughput compared to the baseline HBM-PIM with 16 FP16 multipliers. However, scaling up the number of multipliers makes the SIMD-based design of HBM-PIM inefficient, as every multiplier must be paired with a 32-bit accumulator, leading to substantial register overhead.
To address this, we choose to maintain the same number of output registers as in HBM-PIM, and increase the input bandwidth to match the computational throughput. Fig. 6(c) right shows the proposed PCU microarchitecture containing 16 processing elements (PEs), which share the same set of FP8 activations during a PIM operation. Specifically, the PCU computes a $1\times 4\times 16$ GEMV tile [^2] Through the rest of this paper, we refer a $n\times k\times l$ tile as the dot product between $n\times k$ activations and $k\times l$ weights., with every PE computing a 4-way dot product. The multiplication involves a 6-bit fixed-point multiplier that multiplies the signed ($s$) input mantissa ($m$) with the decoded weight or KV-cache. The multiplier bit-width is determined in the following ways: First, computing the attention-score requires a 5-bit mantissa (including the hidden bit), and the activation requires a sign bit, resulting in a total of 6 bits. Second, since KV-cache adopts INT4-Asym quantization, its precision is 5-bit after adding a 4-bit zero point ($z_{\text{KV}}$). On the other hand, weights are encoded with the BitMoD data type [bitmod], which requires 6 bits in fixed-point representation. After multiplication, the 4-bit input exponent ($e$) is used to shift the product. The results from 4 multipliers are reduced through a 4:2{4:2} compressor tree, followed by 32-bit fixed-point accumulation. Thus, the proposed PE eliminates the area-hungry components such as high-precision multipliers and exponent alignment modules in the conventional FP16/FP32 MAC unit.
@@ V-C P3-LLM Microarchitecture
V-D Throughput-Enhanced PCU Design
Despite the proposed PCU achieving $4\times$ the computational throughput than HBM-PIM, it still supports only GEMV operations. To further exploit data reuse opportunities inside the PCU, one can increase the input size to compute a $n\times 4\times 16$ GEMM tile, where $n>1$. But this will increase the number of multipliers by $n\times$, leading to considerable area overhead. A direct workaround is also to reduce the weight size by $n\times$, resulting in a $n\times 4\times(16/n)$ GEMM tile without the need for additional multipliers. Nevertheless, this approach causes severe under-utilization of the PIM bandwidth because of the inherent 256-bit column access granularity of HBM.
Instead of adopting spatial data reuse as mentioned above, we introduce a novel optimization that exploits temporal data reuse inside the PCU. Our key insight is that the proposed low-precision PCU not only reduces the area, but also permits a higher operating frequency compared to the baseline HBM-PIM [samsung-pim] using FP32 accumulators. Specifically, our evaluation (detailed in Section VI-A) demonstrates that the PCU of P3-LLM can reliably achieve $2\times$ higher frequency than that of HBM-PIM. Given that HBM-PIM operates at a clock cycle of $t_{\text{CCD\_L}}$, which corresponds to 4 memory bus cycles, the proposed PCU can instead operate P3-LLM at $t_{\text{CCD\_S}}$, which is typically a half of $t_{\text{CCD\_L}}$ across all HBM generations [hbm-jedec, hbm3-jedec, hbm4-jedec]. Fig. 7 compares the command timing between HBM-PIM and P3-LLM. Initially, the host NPU sends input activations to the PCU register. Subsequently, a PIM command is initiated every $t_{\text{CCD\_L}}$ on HBM-PIM, which reads out 256-bit weights from the DRAM row buffer and performs a GEMV operation. P3-LLM allows the same weight slice to be reused twice by two different inputs, effectively doubling the throughput for cases where multiple inputs share the same weight (i.e., GQA and non-single-batch linear layers).
@@ VI-A Experimental Methodology
Accelerator Baselines. To evaluate the hardware performance, we compare P3-LLM against three baseline systems: (1) An NPU accelerator running FP16 models, without PIM support in its DRAM; (2) An NPU accelerator integrated with HBM-PIM [samsung-pim] running FP16 models; (3) A SoTA LLM accelerator, Ecco [ecco], employing W4A8KV4 quantization with k-means codebooks and Huffman encoding. All baseline systems share the same NPU and DRAM configurations as P3-LLM with comparable total area. For comparison of accelerator performance, we focus on the decoding stage that dominates the end-to-end inference latency [distserve, zigzag-llm]. We use a context length of 4K, which covers most edge scenarios as suggested in prior works [h2-llm, anda]. Notably, for accelerator evaluation, we only assess Llama-2-(7B, 13B), Llama-3.1-8B, Llama-3.2-3B, and Mistral-7B-v0.3, since the remaining models share the same size and/or architecture.
@@ VI-B Model Performance
Perplexity Results. Table IV summarizes the model perplexity of different quantization methods. For KV-cache quantization, on average, P3-LLM outperforms Oaken and incurs only <0.1{<0.1} perplexity loss. Oaken also has a higher effective KV-cache precision of $4.8$ bits due to its inefficient offline calibration, which keeps a large ratio ($\sim 10\%$}) of outliers in high precision. On the other hand, P3-LLM employs per-head INT4-Asym quantization, where every 128 elements (i.e., the head dimension size of all evaluated models) share a 16-bit scaling factor and a 4-bit zero-point, resulting in an effective precision of $4.16$ bits. Hence, P3-LLM achieves better perplexity even under a lower precision, demonstrating the effectiveness of our dynamic input-aware smoothing for mitigating outliers. For weight-activation quantization, on average, P3-LLM outperforms QuaRot and QServe with $39\%$ and $22\%$ lower perplexity loss, respectively, despite performing more aggressive quantization with 8-bit queries and attention-scores. This highlights the strength of employing our operand-dependent, hybrid numerical formats in reducing quantization error and preserving model quality.
@@
Accuracy Results. Table V shows the model accuracy of different quantization methods. On average, P3-LLM achieves $1.04\%$ higher accuracy than Oaken under 4-bit KV-cache quantization. This improvement is attributed to the dynamic smoothing strategy of P3-LLM, which effectively suppresses the outlier channels in an input-aware manner without overfitting to a specific calibration dataset. Regarding weight-activation quantization, P3-LLM improves the average accuracy by a large margin of $2.57\%$ and $3.05\%$ compared to QuaRot and QServe, respectively. Moreover, the accuracy improvement is particularly pronounced on Llama-3.2-3B. Both QuaRot and QServe use integer formats that do not fit well to LLM tensor distributions, and rely on calibration datasets that lead to severe overfitting, particularly for smaller models with limited knowledge capacity.

### v2

@@ IV-B Attention-Score Quantization
In order to fully exploit the hardware efficiency of KV-cache quantization, it is essential to quantize additional operands, e.g., attention-scores, within the self-attention module. Although attention-scores have little impact on the overall memory footprint (see Section III-A), they play a critical role in the computation flow during self-attention. Consider a low-precision PCU supporting 8-bit multiplication, if the attention scores remain in FP16, the PCU cannot be fully utilized to accelerate self-attention. Instead, the quantized value cache must be transferred to NPU and rely on the FP16 compute units to perform multiplication with attention-scores, thus diminishing the advantages of high internal PIM bandwidth for LLM decoding. Unfortunately, existing KV-cache quantization approaches, such as QServe [qserve] and Oaken [oaken], typically keep attention-scores in FP16, which hinders the usage of low-precision hardware for multiplication with the value cache. A straightforward solution is to quantize attention-scores with INT8 [sage-attn], but this introduces two drawbacks. First, INT8 attention-scores cause noticeable perplexity degradation as shown in Table II. Second, as we will demonstrate in Section IV-D, INT8 can lead to significant accuracy loss for activation quantization. Since attention-scores behave the same as activations during the MAC operation on hardware, it is desirable to identify a good numerical format that balances quantization accuracy and hardware complexity.
To address this, we propose an unsigned 8-bit floating-point format, FP8-S0E4M4, that contains a 4-bit mantissa and a 4-bit exponent with an exponent bias of 15-15. There are two insights that drive our design choice. First, because the attention-scores are produced after a softmax function, its numerical range always lies between 0 and $1$, eliminating the need for a sign bit in its encoding. Second, recall FP16 has a 5-bit exponent with a bias of 15-15, which provides an exponent range of [14,15][-14,15]. However, since attention-scores are always less than 1, there is no need to use positive exponent values, leading to an effective exponent range of [14,1][-14,-1] with 14 distinct values. Thus, a 4-bit exponent is sufficient to represent the attention-score range, which leaves 4 bits to store mantissa with better numerical fidelity. Table II reflects that FP8-S0E4M4 outperforms both INT8 and FP8-E4M3 for attention-score quantization, and achieves near-lossless model performance.
@@ IV-C Weight Quantization
For weight quantization, we build upon an existing numerical format, BitMoD [bitmod], which adaptively remaps the redundant negative zero encoding of FP4 to some pre-defined special values. On top of the basic FP4 quantization values {± $\{\pm\,0,\,\pm\,0.5,\,\pm\,1,\,\pm\,1.5,\,\pm 2,\,\pm 3,\,\pm 4,\,\pm 6\}$, BitMoD introduces four additional special values {± $\{\pm\,5,\,\pm\,8\}$, where one of them can be selected to replace the negative zero. By searching for the optimal special value for each group of weights, it reduces the quantization error compared to asymmetric integer quantization with minimal hardware overhead. Notably, since weights and KV-cache are mapped to the same operand on the MAC hardware, a decoder is required to accommodate the data types of both weights and KV-cache (i.e., BitMoD and INT4-Asym), which we discuss in Section V-A.
@@ V P3-LLM Architecture
This section outlines the architectural details of P3-LLM, which features: (a) a high-throughput and area-efficient PCU co-designed to support hybrid numerical formats, (b) an efficient operator mapping that maps most memory-intensive operators to the low-precision PIM hardware during decoding, (c) a quantized matrix multiplication dataflow that applies operator fusion to minimize the runtime overhead of quantization and dequantization.
V-A P3-LLM Microarchitecture
To support the proposed W4A8KV4P8 scheme with hybrid numerical formats in PIM, we design an efficient PCU on top of HBM-PIM [samsung-pim]. Fig. 6(a) left shows the P3-LLM architecture containing a host NPU and an HBM-PIM. One PIM channel contains 8 PCUs, each of them is shared by 2 banks to amortize the area overhead. During the decoding stage, the NPU sends activations and attention-scores to the input register of PCU, while weights and KV-cache are read from DRAM banks to perform MAC operations. Given that weights and KV-cache are quantized to 4 bits, we choose to equip each PCU with 64 multipliers to match the 256-bit column access granularity of HBM, effectively quadrupling the compute throughput compared to the baseline HBM-PIM with 16 FP16 multipliers. However, scaling up the number of multipliers makes the SIMD-based design of HBM-PIM inefficient, as every multiplier must be paired with a 32-bit accumulator, leading to substantial register overhead.
To address this, we choose to maintain the same number of output registers as in HBM-PIM, and increase the input bandwidth to match the computational throughput. Fig. 6(a) right shows the proposed PCU microarchitecture containing 16 processing elements (PEs), which share the same set of FP8 activations during a PIM operation. Specifically, the PCU computes a $1\times 4\times 16$ GEMV tile [^2] Through the rest of this paper, we refer a $n\times k\times l$ tile as the dot product between $n\times k$ activations and $k\times l$ weights., with every PE computing a 4-way dot product. The multiplication involves a 6-bit fixed-point multiplier that multiplies the signed ($s$) input mantissa ($m$) with the decoded weight or KV-cache. The multiplier bit-width is determined in the following ways: First, computing the attention-score requires a 5-bit mantissa (including the hidden bit), and the activation requires a sign bit, resulting in a total of 6 bits. Second, since KV-cache adopts INT4-Asym quantization, its precision is 5-bit after adding a 4-bit zero point ($z_{\text{KV}}$). On the other hand, weights are encoded with the BitMoD data type [bitmod], which requires 6 bits in fixed-point representation. After multiplication, the 4-bit input exponent ($e$) is used to shift the product. The results from 4 multipliers are reduced through a 4:2{4:2} compressor tree, followed by 32-bit fixed-point accumulation. Thus, the proposed PE eliminates the area-hungry components such as high-precision multipliers and exponent alignment modules in the conventional FP16/FP32 MAC unit.
V-B PIM Offloading Opportunities with P3-LLM
@@
The proposed W4A8KV4P8 quantization allows memory-bound matrix operations to be efficiently executed on the low-precision PIM hardware, as illustrated in Fig. 6(b). During the decoding stage, the host NPU performs all element-wise operations such as RoPE and softmax in high precision, while GEMM/GEMV operators such as linear layers are offloaded to PIM. One exception is $Q\cdot K^{\text{T}}$}, whose mapping depends on whether the key cache is quantized pre-RoPE or post-RoPE. In the former case, the quantized key cache lacks positional information during decoding, which can produce incorrect outputs if multiplied directly with the query. Therefore, an online RoPE operation must be applied to the key cache in each decoding iteration. Fortunately, RoPE is essentially an element-wise operation with negligible runtime overhead[flashinfer]. Since RoPE is performed on NPU and its outputs are high-precision, we opt to calculate $Q\cdot K^{\text{T}}$} on NPU if the key cache is quantized before RoPE. In contrast, the post-RoPE quantized key cache can be directly multiplied with the query, which permits $Q\cdot K^{\text{T}}$} to be computed on low-precision PIM.
V-C Optimization on Quantized Computation Flow
@@
Quantization Granularity. The quantization granularity has a direct impact on the computation flow. For query and KV-cache, we employ per-head quantization, where a group of $D_{h}$ elements are quantized together and $D_{h}$ is the head size. This allows immediate quantization of each newly generated token and eliminates the need to buffer a window of FP16 KV-cache as done in prior works [mant, kivi]. Additionally, weights are quantized per-group with a group size of $128$ following BitMoD [bitmod]. Activations are quantized per-token using FP8-E4M3. Finally, attention-score quantization does not require scaling since FP8-S0E4M4 already covers the required numerical range. Instead, we directly round the least significant bits of every FP16 attention-score and keep its highest 4 mantissa bits.
@@
Quantized Computation Flow. As shown in Fig. 6(c), P3-LLM implements operator fusion to minimize the runtime overhead of (de)quantization, which needs to be executed on NPU using high-precision arithmetic. For linear layers, the dequantization scaling is performed after matrix multiplication. For $Q\cdot K^{\text{T}}$}, since the post-RoPE key cache contains per-channel smoothing factors ($S_{\text{SF}}$), we fuse $S_{\text{SF}}$ into query via element-wise multiplication prior to FP8 quantization. Similarly, for $P\cdot V$}, we fuse the per-value-head scaling factors ($S_{V}$) into attention-scores before FP8 quantization. One potential issue of fusing $S_{V}$ is that it may cause the resulting attention-scores’ magnitude to become larger than one, which is outside the range of FP8-S0E4M4. To prevent this, we apply a second-level scaling that divides $S_{V}$ with its maximum, resulting in $S_{V}\in[0,1]$}. Note that the quantization scaling factor is always positive by definition in Eq. 1, so fusing $S_{V}$ still produces unsigned attention-scores. This second-level scaling factor is later multiplied back into the result of $P\cdot V$}.
@@
V-D Throughput-Enhanced PCU
Despite the proposed low-precision PCU achieving $4\times$ the computational throughput than HBM-PIM, it still supports only GEMV operations. To further exploit data reuse opportunities inside the PCU, one can increase the activation tile size from $1\times 4$ to $n\times 4$ (where $n>1$), and compute a $n\times 4\times 16$ GEMM tile. But this will increase the number of multipliers by $n\times$, leading to considerable area overhead. A direct workaround is also to reduce the weight size by $n\times$, resulting in a $n\times 4\times(16/n)$ GEMM tile without the need for additional multipliers. Nevertheless, this approach causes severe under-utilization of the PIM bandwidth because of the inherent 256-bit column access granularity of HBM.
Instead of adopting spatial data reuse as mentioned above, we introduce a novel optimization that exploits temporal data reuse inside the PCU. Our key insight is that the proposed low-precision PCU not only reduces the area, but also permits a higher operating frequency compared to the baseline HBM-PIM [samsung-pim] that uses FP32 accumulators. Specifically, our evaluation (detailed in Section VI-A) demonstrates that the PCU of P3-LLM can reliably achieve $2\times$ higher frequency than that of HBM-PIM. Given that HBM-PIM operates at a clock cycle of $t_{\text{CCD\_L}}$, which corresponds to 4 memory bus cycles, the proposed PCU can instead operate P3-LLM at $t_{\text{CCD\_S}}$, which is typically a half of $t_{\text{CCD\_L}}$ across all HBM generations [hbm-jedec, hbm3-jedec, hbm4-jedec]. Fig. 7 compares the command timing between HBM-PIM and P3-LLM. Initially, the host NPU sends input activations to the PCU register. Subsequently, a PIM command is initiated every $t_{\text{CCD\_L}}$ on HBM-PIM, which reads out 256-bit weights from the DRAM row buffer and performs a GEMV operation. P3-LLM allows the same weight slice to be reused twice by two different inputs, effectively doubling the throughput for cases where multiple inputs share the same weight (i.e., GQA and non-single-batch linear layers).
@@ VI-A Experimental Methodology
Accelerator Baselines. To evaluate hardware performance, we compare P3-LLM against three baselines: (1) An NPU accelerator running FP16 models, without PIM support in its DRAM; (2) An NPU accelerator integrated with HBM-PIM [samsung-pim] running FP16 models; (3) A SoTA LLM accelerator, Ecco [ecco], employing W4A8KV4 quantization with k-means codebooks and Huffman encoding. All baseline systems share the same NPU and DRAM configurations as P3-LLM with comparable total area. For comparison of accelerator performance, we focus on the decoding stage that dominates the end-to-end inference latency [distserve, zigzag-llm]. We use a context length of 4K, which covers most edge scenarios as suggested in prior works [h2-llm, anda]. Notably, for accelerator evaluation, we only assess Llama-2-(7B, 13B), Llama-3.1-8B, Llama-3.2-3B, and Mistral-7B-v0.3, since the remaining models share the same size and/or architecture.
@@ VI-B Model Performance
Perplexity Results. Table IV summarizes the model perplexity of different quantization methods. For KV-cache quantization, on average, P3-LLM outperforms Oaken and incurs only <0.1{<0.1} perplexity loss. Oaken also has a higher effective KV-cache precision of $4.8$ bits due to its inefficient offline calibration, which keeps a large ratio ($\sim 10\%$}) of outliers in high precision. On the other hand, P3-LLM employs per-head INT4-Asym quantization, where every 128 elements (i.e., the head dimension size of all evaluated models) share a 16-bit scaling factor and a 4-bit zero-point, resulting in an effective precision of $4.16$ bits. Hence, P3-LLM achieves better perplexity even under a lower precision, demonstrating the effectiveness of our dynamic input-aware smoothing for mitigating outliers. For weight-activation quantization, on average, P3-LLM outperforms QuaRot and QoQ with $39\%$ and $22\%$ lower perplexity loss, respectively, despite performing more aggressive quantization with 8-bit queries and attention-scores. This highlights the strength of employing our operand-dependent, hybrid numerical formats in preserving model quality under aggressive quantization.
@@
Accuracy Results. Table V shows the accuracy of different quantization methods. On average, P3-LLM achieves $1.04\%$ higher accuracy than Oaken under 4-bit KV-cache quantization. Regarding weight-activation quantization, P3-LLM improves the average accuracy by a large margin of $2.57\%$ and $3.05\%$ compared to QuaRot and QoQ, respectively. Moreover, the accuracy improvement is particularly pronounced on Llama-3.2-3B. Both QuaRot and QoQ use integer formats that do not fit well to LLM tensor distributions, and rely on calibration datasets that lead to severe overfitting, particularly for smaller models with limited knowledge capacity.